Calculate the wavelength associated with the transition given below. ninitial = 6 and nfinal= 1 Use units of nm.
Answer #1

We will use the Rydberg equation

Where

= wavelength of transition

R = Rydberg constant = 1.0974 * 107 m-1 = 1.0974 * 10-2 nm-1

Given

ninitial =6

nfinal= 1

Substituting the values ,we get

1/ = (1.0974 * 10-2 nm-1) [ (1/ 1)2 - (1/ 6)2]

=0.010669 nm-1

= 1/ (0.010669 nm-1) =93.73 nm

=93.73 nm